OpenStudy (anonymous):

What volume of a 0.25 mol/L salt solution in a lab contains 0.10 mol of sodium chloride?

OpenStudy (anonymous):

molarity= moles x vol

OpenStudy (anonymous):

isnt it 0.4L

OpenStudy (anonymous):

molarity is molarity= mol/l

OpenStudy (accessdenied):

Yes, molarity is equal to the number of moles solute divided by the Liters of solution. We are given the moles solute (0.10 mol NaCl). We know the Molarity (0.25 mol/L). Now we just have to set up an equation to find the volume.

OpenStudy (anonymous):

so 0.25=0.1/y

OpenStudy (anonymous):

y=0.4L?

OpenStudy (accessdenied):

Looks good. :)

OpenStudy (anonymous):

Yeah, it cant be 2.5L right, it doesn't seem right

OpenStudy (accessdenied):

If you checked: 0.10 / 2.5 = 0.04 mol/L, but it isn't equal to 0.25 like we were given.

OpenStudy (anonymous):

True. Thanks for the help

OpenStudy (accessdenied):

You're welcome! :)